X = [0.01, 0.36, 0.97, 0.44, 0.86, 0.49, 0.02, 0.19, 0.54, 0.63, 0.69, 0.27, 0.21, 0.55, 0.54, 0.1 , 0.02, 0.49 , 0.42, 0.79], is a list of 20 random numbers generated from a pseudo-random number generator G.


a. What is a uniformity test?

b. Using Chi-square Goodness-of-fit test with 4 bins, test whether X belongs to U(0, 1) based on 5% significant level. 

c. Can we conclude G generates truly random numbers following the distribution U(0, 1)?


1
Expert's answer
2021-12-10T13:21:45-0500

a.

"\chi^2" squared test for uniformity can be formulated by binning the ranks 0: M into J bins and testing that the bins all have roughly the expected number of draws in them.

If bj is the number of ranks that fall into bin j and ej is the number of ranks expected to fall into

bin j, the test statistic is


"\chi^2=\sum \frac{(b_j-e_j)^2}{e_j}"


b.

"H_0:" X belongs to U(0, 1)

"H_a:" X does not belong to U(0, 1)


"X=[0.01, 0.02, 0.02, 0.1, 0.19, 0.21, 0.27, 0.36, 0.42, 0.44, 0.49, 0.49,"

"0.54, 0.54, 0.55, 0.63, 0.69, 0.79, 0.86, 0.97]"

4 bins:

"[0,0.25],[0.25,0.5],[0.5,0.75],[0.75,1]"


"\chi^2=\sum \frac{(b_j-e_j)^2}{e_j}"


expected number of values in each interval is

"e_j=20\/4=5"

then:


"\chi^2=\frac{(6-5)^2+(6-5)^2+(5-5)^2+(3-5)^2}{5}=6/5=1.2"


"df=4-1=3"


critical value:

"\chi^2_{crit}=0.216"


c.

Since "\chi^2>\chi^2_{crit}" we accept the null hypothesis. X belongs to U(0, 1)